# Adapting Multilingual LLMs to Low-Resource Languages using Continued Pre-training and Synthetic Corpus

### v1

@@ Abstract
Multilingual LLMs support a variety of languages; however, their performance is suboptimal for low-resource languages. In this work, we emphasize the importance of continued pre-training of multilingual LLMs and the use of translation-based synthetic pre-training corpora for improving LLMs in low-resource languages. We conduct our study in the context of the low-resource Indic language Hindi. We introduce Nemotron-Mini-Hindi 4B, a bilingual SLM supporting both Hindi and English, based on Nemotron-Mini 4B. The model is trained using a mix of real and synthetic Hindi + English tokens, with continuous pre-training performed on 400B tokens. We demonstrate that both the base and instruct models achieve state-of-the-art results on Hindi benchmarks while remaining competitive on English tasks. Additionally, we observe that the continued pre-training approach enhances the model’s overall factual accuracy.
@@ 1 Introduction
Recently, in the context of Indic languages, target language Supervised Fine-Tuning (SFT) has become a common practice to adapt LLMs to specific languages Gala et al. (2024). However, it remains to be studied whether language-specific tuning improves LLMs’ understanding in regional contexts. Some studies suggest that SFT can introduce LLMs to new domain knowledge, though it is typically used to enhance the model’s instruction-following capability Mecklenburg et al. (2024). SFT on translated English instruction tuning data is widely used to develop regional LLMs for Indic languages. While this may improve instruction-following in the target language, it may not enhance LLMs’ understanding of regional contexts Balachandran (2023). Another approach to updating LLM knowledge is continued pre-training, but the limited availability of tokens for low-resource languages makes this both infeasible and prone to overfitting.
@@
Based on this approach, we present Nemotron-Mini-Hindi-4B-Base [^1] https://huggingface.co/nvidia/Nemotron-4-Mini-Hindi-4B-Base and Nemotron-Mini-Hindi-4B-Instruct [^2] https://huggingface.co/nvidia/Nemotron-4-Mini-Hindi-4B-Instruct [^3] https://build.nvidia.com/nvidia/nemotron-4-mini-hindi-4b-instruct, state-of-the-art SLMs for the Hindi language. These SLMs support Hindi, English, and Hinglish. The Hindi models are based on the multilingual Nemotron-Mini-4B (also known as Minitron-4B), adapted with continued pre-training on 400 billion Hindi and English tokens. The data blend used equal proportions of both languages. The instruct version of the model was developed using SFT and DPO techniques. The model outperforms all similarly sized models on various IndicXTREME, IndicNLG benchmark tasks and popular translated English benchmarks such as MMLU, Hellaswag, ARC-C, and ARC-E Gala et al. (2024). We also perform LLM-based evaluations using the benchmark datasets IndicQuest Rohera et al. (2024) and in-house SubjectiveEval, with GPT-4 serving as the judge LLM. This is the first study to present and evaluate bilingual language models of this nature. We provide a thorough study of the models in both languages.
@@ 2 Related Work
Cahyawijaya et al. (2024) shows that large language models can learn low-resource languages effectively using in-context learning and few-shot examples, improving performance through cross-lingual contexts without extensive tuning. Gurgurov et al. (2024) enhances multilingual LLMs for low-resource languages by using adapters with data from ConceptNet, boosting performance in sentiment analysis and named entity recognition.
@@ 3 Methodology
In this section, we describe our methodology for adapting multilingual LLMs to target languages to improve performance in those languages. Specifically, we build a bilingual SLM that supports both Hindi and English. We conduct our adaptation experiments using the multilingual Nemotron-Mini-4B model (also known as Minitron-4B). The model undergoes continuous pre-training with an equal mixture of Hindi and English data, consisting of 200B tokens per language. The original Nemotron-4B model was primarily trained on English tokens and had seen only 20B Hindi tokens. Given the limited amount of Hindi data, adapting an existing multilingual model rather than training from scratch is an effective strategy, allowing us to leverage the knowledge learned from the pre-trained model. Additionally, as Nemotron-4B employs a large 256k tokenizer, we did not need to extend the tokenizer. The fertility ratio for Hindi text is 1.7, which is better than that of its Llama and Gemma counterparts.
@@ 3.1 Synthetic Data Curation
The translated Hindi data comprises approximately 60 billion tokens. We then combine this synthetic data with around 40 billion real tokens (web-scraped data) to create a dataset totaling 100 billion Hindi tokens. Additionally, this entire Hindi text is transliterated into Roman script, expanding the total dataset to 220 billion tokens. The transliterated tokens are included to enable the model to support Hinglish queries. This Hindi data is further combined with 200 billion English tokens for continued pre-training. Including the English dataset helps prevent catastrophic forgetting of English capabilities and contributes to training stability. Fuzzy deduplication is performed on the entire text using NeMo-Curator [^6] https://github.com/NVIDIA/NeMo-Curator to eliminate similar documents. The real Hindi data sources include internal web-based datasets and Sangraha Corpus Singh et al. (2024). The English dataset is a subset of the pre-training corpus used for the Nemotron-15B model. All the datasets used in this work are commercially friendly.
@@ 3.2 Continued Pre-training
Base models Metric Nemotron-Mini-Hindi-4B Nemotron-Mini-4B Sarvam 2B Gemma 2-2B Openhathi Llama-3.1 8B Gemma 2-9B
IndicSentiment F1 - NLU 84.31 72.47 81.90 91.90 72.89 92.06 94.90
IndicCopa F1 - NLU 81.86 62.50 64.08 58.65 68.69 61.87 72.58
IndicXNLI F1 - NLU 49.67 40.39 28.53 16.67 16.67 16.67 16.79
IndicXParaphrase F1 - NLU 37.09 16.27 25.29 26.60 71.72 72.75 71.38
Indic QA (With Context) 1 shot F1 - NLG 18.32 15.10 3.90 33.37 20.69 35.92 46.27
Indic Headline 1 shot BLEURT - NLG 0.50 0.46 0.21 0.27 0.47 0.38 0.27
IndicWikiBio 1 shot BLEURT - NLG 0.62 0.59 0.56 0.60 0.52 0.60 0.63
MMLU Acc - NLU 49.89 38.20 26.75 35.05 32.27 44.84 55.08
BoolQ Acc - NLU 71.71 70.79 40.00 66.00 58.56 61.00 61.00
ARC Easy Acc - NLU 78.81 58.25 32.37 52.31 44.28 67.05 85.69
Arc Challenge Acc - NLU 65.02 47.87 27.81 40.78 32.68 54.10 76.02
Hella Swag Acc - NLU 31.66 25.31 24.70 27.50 25.59 33.50 42.40
@@ 3.3 Model Alignment
The first alignment stage is Supervised Fine-Tuning (SFT). We use a general SFT corpus with approximately 200k examples, comprising various tasks as outlined in Adler et al. (2024). The model is trained for one epoch with a global batch size of 1024 and a learning rate in the range of [5e-6, 9e-7], using cosine annealing. Due to the lack of a high-quality Hindi SFT corpus, we leverage English-only data for SFT. We also experimented with translated English data (filtered using back-translation-based methods) for SFT, but did not observe any improvements with this addition. We found that using the English-only SFT corpus enhances instruction-following capabilities in Hindi, highlighting the cross-lingual transferability of these skills.
@@ 6 Limitations
To mitigate these issues to some extent, we have implemented safety alignment during the DPO stage to guide the model away from responding to toxic or harmful content. Additionally, we conduct safety evaluations using benchmarks such as Aegis [^7] https://huggingface.co/datasets/nvidia/Aegis-AI-Content-Safety-Dataset-1.0, Garak [^8] https://github.com/leondz/garak, and Human Content red-teaming, and our findings indicate that the model’s responses remain within permissible limits.
@@ Acknowledgements
This work would not have been possible without contributions from many people at NVIDIA. To mention a few: Asif Ahamed, Ayush Dattagupta, Umair Ahmed, Yoshi Suhara, Ameya Mahabaleshwarkar, Zijia Chen, Varun Singh, Vibhu Jawa, Rohit Watve, Oluwatobi Olabiyi, Mostofa Patwary, and Oleksii Kuchaiev.

### v2

@@ Abstract
Multilingual LLMs support a variety of languages; however, their performance is suboptimal for low-resource languages. In this work, we emphasize the importance of continued pre-training of multilingual LLMs and the use of translation-based synthetic pre-training corpora for improving LLMs in low-resource languages. We conduct our study in the context of the low-resource Indic language Hindi. We introduce Nemotron-Mini-Hindi 4B, a bilingual SLM supporting both Hindi and English, based on Nemotron-Mini 4B. The model is trained using a mix of real and synthetic Hindi + English tokens, with continuous pre-training performed on 400B tokens. We demonstrate that both the base and instruct models achieve state-of-the-art results on Hindi benchmarks while remaining competitive on English tasks. Additionally, we observe that the continued pre-training approach enhances the model’s overall factual accuracy. We perform an ablation study to highlight the impact of Hindi pre-training, showing significant improvements in Hindi chat capabilities and factual accuracy, which cannot be achieved through Hindi alignment alone.
@@ 1 Introduction
Recently, in the context of Indic languages, target language Supervised Fine-Tuning (SFT) has become a common practice to adapt LLMs to specific languages Gala et al. (2024). However, it remains to be studied whether language-specific SFT tuning improves LLMs’ understanding in regional contexts. Some studies suggest that SFT can introduce LLMs to new domain knowledge, though it is typically used to enhance the model’s instruction-following capability Mecklenburg et al. (2024). SFT on translated English instruction tuning data is widely used to develop regional LLMs for Indic languages. While this may improve instruction-following in the target language, it may not enhance LLMs’ understanding of regional contexts Balachandran (2023). Another approach to updating LLM knowledge is continued pre-training, but the limited availability of tokens for low-resource languages makes this both infeasible and prone to overfitting.
@@
Based on this approach, we present Nemotron-Mini-Hindi-4B-Base [^1] https://huggingface.co/nvidia/Nemotron-4-Mini-Hindi-4B-Base and Nemotron-Mini-Hindi-4B-Instruct [^2] https://huggingface.co/nvidia/Nemotron-4-Mini-Hindi-4B-Instruct [^3] https://build.nvidia.com/nvidia/nemotron-4-mini-hindi-4b-instruct, state-of-the-art Small Language Models (SLMs) for the Hindi language. These SLMs support Hindi, English, and Hinglish. The Hindi models are based on the multilingual Nemotron-Mini-4B (also known as Minitron-4B), adapted with continued pre-training on 400 billion Hindi and English tokens. The data blend used equal proportions of both languages. The instruct version of the model was developed using SFT and DPO techniques. The model outperforms all similarly sized models on various IndicXTREME, IndicNLG benchmark tasks and popular translated English benchmarks such as MMLU, Hellaswag, ARC-C, and ARC-E Gala et al. (2024). We also perform LLM-based evaluations using the benchmark datasets IndicQuest Rohera et al. (2024) and in-house SubjectiveEval, with GPT-4 serving as the judge LLM. This is the first study to present and evaluate bilingual language models of this nature. We provide a thorough study of the models in both languages.
Additionally, we perform an extensive ablation to analyze the impact of Hindi SFT and DPO on both the base multilingual Nemotron-Mini-4B and the Hindi-specific Nemotron-Mini-Hindi-4B. Our results show that while Hindi alignment improves performance, Hindi pre-training is essential for achieving strong results on Hindi benchmarks. Notably, Nemotron-Mini-Hindi 4B leads to significant gains in factual accuracy, not just for Hindi but also for English, showcasing effective cross-lingual transfer.
@@ 2 Related Work
Cahyawijaya et al. (2024) show that large language models can learn low-resource languages effectively using in-context learning and few-shot examples, improving performance through cross-lingual contexts without extensive tuning. Gurgurov et al. (2024) enhance multilingual LLMs for low-resource languages by using adapters with data from ConceptNet, boosting performance in sentiment analysis and named entity recognition.
@@ 3 Methodology
In this section, we describe our methodology for adapting multilingual LLMs to target languages to improve performance in those languages. Specifically, we build a bilingual SLM that supports both Hindi and English. We conduct our adaptation experiments using the multilingual Nemotron-Mini-4B model (also known as Minitron-4B). The model undergoes continuous pre-training with an equal mixture of Hindi and English data, consisting of 200B tokens per language. The original Nemotron-4B model was primarily trained on English tokens and had seen only 20B Hindi tokens. Given the limited amount of Hindi data, adapting an existing multilingual model rather than training from scratch is an effective strategy, allowing us to leverage the knowledge learned from the pre-trained model. Additionally, as Nemotron-4B employs a large 256k tokenizer, we did not need to extend the tokenizer. The fertility ratio for Hindi text is 1.7, which is better than that of its Llama (2.64) and Gemma (1.98) counterparts.
@@ 3.1 Synthetic Data Curation
The translated Hindi data comprises approximately 60 billion tokens. We then combine this synthetic data with around 40 billion real tokens (web-scraped data) to create a dataset totaling 100 billion Hindi tokens. Additionally, this entire Hindi text is transliterated into Roman script, expanding the total dataset to 220 billion tokens. The transliterated tokens are included to enable the model to support Hinglish queries. This Hindi data is further combined with 200 billion English tokens for continued pre-training. Including the English dataset helps prevent catastrophic forgetting of English capabilities and contributes to training stability. Fuzzy deduplication is performed on the entire text using NeMo-Curator [^6] https://github.com/NVIDIA/NeMo-Curator to eliminate similar documents. The real Hindi data sources include internal web-based datasets and Sangraha Corpus Khan et al. (2024). The English dataset is a subset of the pre-training corpus used for the Nemotron-15B model. All the datasets used in this work are commercially friendly.
@@ 3.2 Continued Pre-training
Base models Metric Nemotron-Mini-Hindi-4B Nemotron-Mini-4B Sarvam-1 2B Gemma 2-2B Openhathi Llama-3.1 8B Gemma 2-9B
IndicSentiment F1 - NLU 84.31 72.47 96.36 91.90 72.89 92.06 94.90
IndicCopa F1 - NLU 81.86 62.50 51.63 58.65 68.69 61.87 72.58
IndicXNLI F1 - NLU 49.67 40.39 36.08 16.67 16.67 16.67 16.79
IndicXParaphrase F1 - NLU 37.09 16.27 80.99 26.60 71.72 72.75 71.38
Indic QA (With Context) 1 shot F1 - NLG 18.32 15.10 35.81 33.37 20.69 35.92 46.27
Indic Headline 1 shot BLEURT - NLG 0.50 0.46 0.36 0.27 0.47 0.38 0.27
IndicWikiBio 1 shot BLEURT - NLG 0.62 0.59 0.53 0.60 0.52 0.60 0.63
MMLU Acc - NLU 49.89 38.20 45.65 35.05 32.27 44.84 55.08
BoolQ Acc - NLU 71.71 70.79 56.08 66.00 58.56 61.00 61.00
ARC Easy Acc - NLU 78.81 58.25 76.85 52.31 44.28 67.05 85.69
Arc Challenge Acc - NLU 65.02 47.87 59.04 40.78 32.68 54.10 76.02
Hella Swag Acc - NLU 31.66 25.31 37.13 27.50 25.59 33.50 42.40
@@
Model Setting SubjectiveEval IndicQuest (Hi) IndicQuest (En)
Nemotron-Mini-4B-Base
   SFT (En) + DPO (En) 1.92 2.66 3.89
   SFT (En) + DPO (En + Hi) 1.88 2.80 3.87
   SFT (En + Hi) + DPO (En) 2.73 3.20 3.86
   SFT (En + Hi) + DPO (En + Hi) 2.51 3.14 3.88
Nemotron-Mini-Hindi-4B-Base
   SFT (En) + DPO (En) 3.81 4.12 4.02
   SFT (En) + DPO (En + Hi) 4.3 4.10 4.03
   SFT (En + Hi) + DPO (En) 4.28 4.06 4.02
   SFT (En + Hi) + DPO (En + Hi) 4.25 4.13 4.04
Table 5: Ablation study of post-training configurations analyzing the impact of Hindi pretraining on SubjectiveEval (Chat capability) and IndicQuest (Factual accuracy) tasks.
@@ 3.3 Model Alignment
The first alignment stage is Supervised Fine-Tuning (SFT). We use a general SFT corpus with approximately 200k examples, comprising various tasks as outlined in Adler et al. (2024). The model is trained for one epoch with a global batch size of 1024 and a learning rate in the range of [5e-6, 9e-7], using cosine annealing. Due to the lack of a high-quality Hindi SFT corpus, we leverage English-only data for SFT. We also experimented with translated English data (filtered using back-translation-based methods) for SFT, but did not observe any improvements with this addition. We found that using the English-only SFT corpus enhances instruction-following capabilities in Hindi, highlighting the cross-lingual transferability of these skills. For the ablation study, we use approximately 70k high-quality Hindi examples selected from a larger pool of 200k translated samples. These Hindi instances are derived by translating the original English data, with noisy or low-quality translations filtered out using a back-translation-based filtering approach.
@@ 4 Results and Discussion
Table 5 shows an ablation study analyzing the impact of Hindi SFT and DPO on Nemotron-Mini-4B Base and the Hindi-pretrained Nemotron-Mini-Hindi-4B Base. While Hindi alignment improves the base Nemotron-Mini-4B, its performance remains significantly lower than the Hindi-pretrained model across all metrics. Notably, Nemotron-Mini-Hindi-4B with English alignment alone outperforms Nemotron-Mini-4B even with Hindi alignment, highlighting the strong benefits of Hindi pretraining. The best results are achieved with English SFT and (English + Hindi) DPO on the Hindi-pretrained model, suggesting that mixed-language alignment can enhance generalization. Overall, Hindi pretraining leads to substantial improvements not only in Hindi understanding but also in factual accuracy and English performance, demonstrating effective cross-lingual transfer.
@@ 5 Conclusion
To mitigate these issues to some extent, we have implemented safety alignment during the DPO stage to guide the model away from responding to toxic or harmful content. Additionally, we conduct safety evaluations using benchmarks such as Aegis [^7] https://huggingface.co/datasets/nvidia/Aegis-AI-Content-Safety-Dataset-1.0 Ghosh et al. (2024), Garak [^8] https://github.com/leondz/garak Derczynski et al. (2024), and Human Content red-teaming, and our findings indicate that the model’s responses remain within permissible limits.
@@ Acknowledgements
This work would not have been possible without contributions from many people at NVIDIA. To mention a few: Asif Ahamed, Ayush Dattagupta, Umair Ahmed, Yoshi Suhara, Ameya Mahabaleshwarkar, Zijia Chen, Varun Singh, Vibhu Jawa, Saurav Muralidharan, Sharath Turuvekere Sreenivas, Marcin Chochowski, Rohit Watve, Oluwatobi Olabiyi, Mostofa Patwary, and Oleksii Kuchaiev.